# Associative Transformer

### v3

@@ 4.6 Prior Specialization via Bottleneck Attention Balance Loss
We aim to investigate whether the bottleneck attention could contribute to specialized priors that guide attention to different localized features. To this end, for each prior in the explicit memory, we visualized its attention scores computed by Equation 2 over different patches in Figure 4, where the selected patches are highlighted. Interestingly, despite employing the squash layer to concatenate patches across different batch samples, each prior eventually learned to extract from similar patch input positions for different samples. These priors demonstrated emergent specialization, focusing on specific spatial areas in an image.
Figure 4: Prior-guided attention visualization highlights selected patches, with each prior having learned to extract from similar patch input positions across different samples. These priors exhibit emergent specialization, focusing on specific spatial areas in an image (using the first global workspace layer of AiT-Small in CIFAR-10).
4.7 Relational Reasoning
Table 5: Relational reasoning task performance.
Methods Relational Non-relational(%)
@@ 4.7 Relational Reasoning
Coordination (Goyal et al., 2022) 73.43 96.31
Set Transformer (Lee et al., 2019) 47.63 57.65
@@
The specialization of priors in the spatial allocation of attention contributes to AiT’s enhanced performance in relational reasoning tasks. The goal is to answer questions about the properties and relations of various objects based on a given image. AiT leverages different priors to facilitate associations among different image regions for the question-answering tasks. In particular, we evaluated AiT’s performance in the Sort-of-CLEVR dataset (Santoro et al., 2017), comprising non-relational and relational tasks. In the non-relational tasks, the question involves the attributes of a specific object, whereas in the relational tasks, the question focuses on the relations between different objects.
To tackle the Sort-of-CLEVR tasks, we embedded the question using a learnable linear projection, with layer normalization applied before and after the embedding. Then, the question embedding was concatenated to the image patch embeddings as the input to AiT. We compared AiT to both Transformer-based models including Set Transformer and Coordination, and other non-Transformer based models including CNN+LSTM (Santoro et al., 2017) and Dense networks. In CNN+LSTM, images are processed with a CNN to produce a set of objects, and questions are processed with an LSTM to generate a question embedding. Then, an MLP combines both the image and question embeddings to answer the question. In addition, Dense-Small and Dense-Base are two fully-connected models derived from AiT-Small and AiT-Base, respectively.
The results in Table 5 demonstrate a substantial improvement in AiT’s performance in both relational and non-relational tasks compared to existing Transformer-based methods of Coordination and Set Transformer. AiT achieved a competitive test accuracy of 80.03% in the relational reasoning tasks. The superior performance with associative memory can be attributed to the inherent sparse attention guided by the explicit memory, allowing the allocation of attention to specific areas of images for different tasks.
5 Limitations and Conclusions
We proposed the Associative Transformer (AiT) with explicit memory and Hopfield networks to enhance associations among sparsely attended input patches. AiT demonstrated largely improved parameter efficiency and performance across various classification and relational reasoning tasks, surpassing Vision Transformers and conventional sparse Transformers. AiT also established a new SOTA performance in the Sort-of-CLEVR dataset, outperforming the Coordination method.
While AiT demonstrates considerable capability, scaling it to larger dataset training without fine-tuning hyperparameters such as the inverse temperature and bottleneck size, remains a challenge. Overall, we are excited about introducing localized contextual learning and associative memory into Transformers by focusing on diverse rather than narrow patch representation learning and selection for tasks involving classification and relational reasoning.
@@ Appendix A Datasets
In this section, we describe the datasets used in this work. (1) CIFAR-10 is an image collection of 10 objects, covering 50k training samples and 10k test samples, labeled as airplane, automobile, and so on. The size of images is $32\times 32\times 3$. (2) CIFAR-100 (Krizhevsky & Hinton, 2009) contains 100 object classes with 500 training images and 100 testing images per class. For both the CIFAR-10 and CIFAR-100 datasets, we performed random cropping with size $32\times 32\times 3$ and a padding size of 4. (3) Triangle dataset (Goyal et al., 2022) includes 50k training images and 10k test images with size $64\times 64$, each of which contains 3 randomly placed clusters of points. The task is to predict whether the three clusters form an equilateral triangle or not. (4) Oxford-IIIT Pet dataset (Parkhi et al., 2012) comprises 37 categories featuring diverse breeds of cats and dogs, with 200 images allocated for each class. We utilized random resized cropping with size $256\times 256\times 3$ and resized all images to size $224\times 224\times 3$. Additionally, we applied random horizontal flip and normalization to the CIFAR-10, CIFAR-100, and Oxford-IIIT Pet datasets. (5) Sort-of-CLEVR dataset (Santoro et al., 2017) is a simplified version of the CLEVR dataset (Johnson et al., 2017). It includes 10k images with size $75\times 75\times 3$ and 20 different questions (10 relational and 10 non-relational questions) for each image. In each image, objects with randomly chosen shapes (square or circle) and randomly chosen colors (red, green, blue, orange, gray, yellow) are placed (Figure 5).
Figure 5: Examples from the Sort-of-CLEVR dataset (Santoro et al., 2017).
@@ Appendix B Comparison of Work Related to Global Workspace Theory
Vision Transformer (Dosovitskiy et al., 2020) $\times$ $\times$ $\checkmark$ $\times$ BU $\times$
BlockBERT (Qiu et al., 2020) $\times$ $\times$ $\checkmark$ $\times$ BU $\checkmark$
BRIMs(Mittal et al., 2020) $\times$ $\times$ $\times$ $\times$ TD $\checkmark$
Modern Hopfield (Ramsauer et al., 2021) $\times$ $\checkmark$ $\times$ $\checkmark$ BU $\times$
Perceiver (Jaegle et al., 2021) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ BU $\times$
Coordination (Goyal et al., 2022) $\checkmark$ $\checkmark$ $\times$ $\times$ BU $\checkmark$
Perceiver IO (Jaegle et al., 2022) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ TD $\times$
Set Transformer (Lee et al., 2019) $\checkmark$ $\checkmark$ $\times$ $\times$ BU $\times$
Luna (Ma et al., 2021) $\checkmark$ $\checkmark$ $\times$ $\times$ BU $\times$
GMAT (Gupta & Berant, 2020) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ BU $\checkmark$
@@ Appendix C Experimental Settings and Hyperparameters
Table 7 presents the hyperparameters used for the different tasks in this study. Unless otherwise noted, we employed the author-recommended settings and hyperparameters for the re-implementation of baseline models. The small variants of ViT and AiT have 2 attention layers, and the base variants of them have 12 attention layers instead. We used the same dimension of the hidden layer and the MLP layer for ViT and AiT. By default, we employed 8 attention heads and 32 memory slots for the bottleneck attention. To obtain the bottleneck size, we considered two main factors of the batch size and the patch size. For the CIFAR and Pet datasets, we used a bottleneck size of 512, which selected from a pool of 32.8k/25.1k patches. For the Triangle dataset, we used a bottleneck size of 64 from a pool of 2.0k patches. For the relational reasoning tasks, we used a bottleneck size of 256, which selected from a pool of 14.4k patches. Based on the bottleneck size and the patch pool size, we used 128 memory slots for the Pet dataset and 32 memory slots for the other datasets. Moreover, we trained the models on the Pet dataset for 300 epochs and on the other datasets for 100 epochs. In relational reasoning tasks, we trained all models for 100 epochs with a batch size of 64.
Table 7: Hyperparameters.
@@
Optimizer
AdamW
Weight decay
0.01
Learning rate
$1\times 10^{-4}$
Number of self-attention heads
12
Number of attention layers
2 (Small)/ 12 (Base)
Size of hidden layer
768
Size of MLP
3072
Size of memory slot
32
Number of bottleneck attention heads
8
1.0
Epochs
100 (300 for Oxford Pet)
@@
Patch size
4
Batch size
512
Number of memory slots
32
Bottleneck size
512
@@
Patch size
32
Batch size
512
Number of memory slots
32
Bottleneck size
64
Oxford Pet
Patch size
16
Batch size
128
Number of memory slots
128
Bottleneck size
512
@@
Patch size
5
Batch size
64
Number of memory slots
32
Bottleneck size
256
Appendix D Analysis of Attention Head Operating Modes
@@
We assume that the competition within the pair-wise attention is important for the model to learn meaningful representations. If such competition exists, a trained model will naturally result in sparser interactions in attention heads. Therefore, we first performed an analysis of the operating modes of different attention heads in a pretrained ViT model by measuring the number of patches each head is attending to. We trained a ViT-Base variant model for 100 epochs from scratch for the CIFAR-10 task. Then, for each attention head, we obtained a violin plot to represent the distribution of attention sparsity for different patches (Figure 6). The attention sparsity for a specific patch’s interactions with other patches is computed as follows $\arg\min_{s}\sum_{j=1}^{s}A^{i,j}\geq 0.9$, where $A^{i,j}$ is the attention score allocated to the $j$th patch by the $i$th patch. The attention sparsity score is measured by the minimal number of required patches whose attention scores add up to 0.90. For instance, there are 65 patches for the CIFAR-10 task with patch size 4, thus there are 65 interactions for each patch to all the patches including itself. An attention head has a higher sparsity if the median $\bar{s}$ of the required patches is smaller, which is the number in the center of each panel. The heads in each layer are sorted according to $\bar{s}$. Note that training the model for a longer duration can result in even better convergence and higher attention sparsity. We also refer to a concurrent investigation on the Bidirectional Encoder Representations from Transformers (BERT) for results on the natural language processing (NLP) tasks (Ramsauer et al., 2021).
@@
Figure 6: Analysis of operating modes of attention heads in the ViT-Base model. We recognize three different groups of attention heads based on their sparsity scores. Group (I) in light blue: High sparsity heads abundant in the middle layers 3-6. The vast majority of these heads only used 50% or fewer interactions. Group (II) in orange: Middle sparsity heads predominant in layers 2 and 7-10. Less than 80% of the interactions were activated. Group (III) in red: Low sparsity heads observed in high layers 11-12 and the first layer, where the most patches were attended to. The global workspace layer will provide the inductive bias to attend to the essential patches more effectively.
@@
Compared to NLP tasks, much less sparsity was found in ViT models. This can be attributed to the long-range dependencies of the attention mechanism were more frequently employed in handling patches of images, compared to handling tokens in text sequences. In ViT models, an input image is divided into non-overlapping patches which are then treated as separate ”tokens” and fed into the transformer model. By contrast, the sequential nature of the NLP text data allows the model to capture long-range dependencies among long-range tokens more easily.
Appendix E Efficacy of the Bottleneck Attention Balance Loss

### v4

@@ 4.6 Prior specialization
The bottleneck attention balance loss encourages competition facilitating the selection of diverse tokens through the bottleneck. We aim to investigate whether the bottleneck attention could contribute to specialized priors that guide attention to various localized features for long-range association. To quantitatively measure the diversity of the selected tokens, we computed the ratio of distinct instances in the selected tokens. Notably, for each prior in the explicit memory, we visualized its attention scores computed by Equation 2 over the different input token positions. Figure 7 presents the corresponding image patches for the selected tokens, with an apparent increase in the patches’ diversity as training progresses. Additionally, Figure 8 highlights these attended patches in the image samples where the learned priors demonstrated emergent spatial specialization, dividing labor by focusing on various aspects of the images.
4.7 Relational reasoning tasks
The specialization in the spatial attention allocation of priors results in AiT’s enhanced performance in relational reasoning tasks. The aim is to answer questions about the properties and relations of various objects in a given image. We evaluated AiT’s performance in the Sort-of-CLEVR dataset [14] for the relational reasoning tasks, which comprises both non-relational and relational tasks. In the non-relational tasks, the question involves the attributes of a specific object, whereas in the relational tasks, the question focuses on the relations between different objects.
Methods Relational (%) Non-relational (%)
@@ 4.7 Relational reasoning tasks
Coordination [13] 73.43 96.31
ViT-Base 63.35 99.73
ViT-Medium 54.71 99.70
ViT-Small 51.75 98.80
Set Transformer [23] 47.63 57.65
@@
Table 6: Relational reasoning tasks.
To tackle the Sort-of-CLEVR task, we embedded a question using a learnable linear projection, with layer normalization employed before and after the embedding. Then, the question embedding was concatenated to the image patch embeddings as the input to AiT. We compared to both Transformer-based models including Set Transformer and Coordination, and other non-Transformer based models including CNN+LSTM [14] and Dense networks. In CNN+LSTM, images were processed with a CNN to produce a set of objects, and questions were processed with an LSTM to generate a question embedding. Then, an MLP combined the image and question embeddings to predict the answer of the question. Dense-Small and Dense-Base are two fully-connected models derived from AiT-Small and AiT-Base. Table 6 demonstrates AiT’s superior performance in the relational and non-relational tasks compared to the Coordination and Set Transformer, achieving a competitive accuracy of 80.03% in the relational tasks. In addition, AiT outperformed the non-Transformer CNN+LSTM model, which relied on a stronger assumption of the input data, i.e., employing the CNN and LSTM for the images and questions, respectively. By contrast, AiT utilized the concatenated tokens of images and questions for the same operations in the proposed Global Workspace Layer.
5 Conclusions
We proposed the Associative Transformer (AiT) based on the bottleneck attention in the explicit memory and the token reconstruction within associative memory to enhance long-range associations of sparsely attended tokens in sparse Transformers. The extensive experiment results demonstrated AiT’s enhanced performance and parameter efficiency in the various classification and relational reasoning tasks, surpassing the Vision Transformers and conventional sparse Transformers. We are excited about inducing associative memory for localized contextual learning in sparse Transformers through the set of diverse priors and the continuous Hopfield networks for various tasks.
Limitations
Scaling AiT to larger dataset training without fine-tuning the Global Workspace Layer’s hyperparameters such as the Hopfield inverse temperature remains a challenge. Implementing a mechanism to adjust adaptively could enhance model performance as it scales. We aim to investigate the countermeasure for more complex tasks, including the CLEVER [40] and bAbI [41] datasets in the future. AiT enables more accessible and cost-effective Transformer model training, showing promise for applications in other domains, such as audio and video. However, potential risks include susceptibility to targeted adversarial attacks, such as Trojan samples exploiting the sparse bottleneck attention mechanism. Exploring potential countermeasures is crucial for its safe deployment in real-world applications.
Acknowledgment
This work was supported in part by Microsoft Research Asia D-CORE Collaborative Research Program.
@@ Appendix A Datasets
In this section, we describe the datasets used in this work. (1) CIFAR-10 is an image collection of 10 objects, covering 50k training samples and 10k test samples, labeled as airplane, automobile, and so on. The size of images is $32\times 32\times 3$. (2) CIFAR-100 [36] contains 100 object classes with 500 training images and 100 testing images per class. For both the CIFAR-10 and CIFAR-100 datasets, we performed random cropping with size $32\times 32\times 3$ and a padding size of 4. (3) Triangle dataset [13] includes 50k training images and 10k test images with size $64\times 64$, each of which contains 3 randomly placed clusters of points. The task is to predict whether the three clusters form an equilateral triangle or not. (4) Oxford-IIIT Pet dataset [37] comprises 37 categories featuring diverse breeds of cats and dogs, with 200 images allocated for each class. We utilized random resized cropping with size $256\times 256\times 3$ and resized all images to size $224\times 224\times 3$. Additionally, we applied random horizontal flip and normalization to the CIFAR-10, CIFAR-100, and Oxford-IIIT Pet datasets. (5) Sort-of-CLEVR dataset [14] is a simplified version of the CLEVR dataset [40]. It includes 10k images with size $75\times 75\times 3$ and 20 different questions (10 relational and 10 non-relational questions) for each image. In each image, objects with randomly chosen shapes (square or circle) and randomly chosen colors (red, green, blue, orange, gray, yellow) are placed (Figure 4).
Figure 4: Examples from the Sort-of-CLEVR dataset [14].
@@ Appendix B Comparison of work related to Global Workspace Theory
Vision Transformer [2] $\times$ $\times$ $\checkmark$ $\times$ BU $\times$
BlockBERT [28] $\times$ $\times$ $\checkmark$ $\times$ BU $\checkmark$
BRIMs[39] $\times$ $\times$ $\times$ $\times$ TD $\checkmark$
Modern Hopfield [22] $\times$ $\checkmark$ $\times$ $\checkmark$ BU $\times$
Perceiver [26] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ BU $\times$
Coordination [13] $\checkmark$ $\checkmark$ $\times$ $\times$ BU $\checkmark$
Perceiver IO [27] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ TD $\times$
Set Transformer [23] $\checkmark$ $\checkmark$ $\times$ $\times$ BU $\times$
Luna [25] $\checkmark$ $\checkmark$ $\times$ $\times$ BU $\times$
GMAT [24] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ BU $\checkmark$
@@ Appendix C Experimental settings and hyperparameters
Table 8 presents the hyperparameters used for the different tasks in this study. Unless otherwise noted, we employed the author-recommended settings and hyperparameters for the re-implementation of baseline models. The small variants of ViT and AiT have 2 attention layers, and the base variants of them have 12 attention layers instead. We used the same dimension of the hidden layer and the MLP layer for ViT and AiT. By default, we employed 8 attention heads and 32 memory slots for the bottleneck attention. To obtain the bottleneck size, we considered two main factors of the batch size and the patch size. For the CIFAR, Pet, and ImageNet100 datasets, we used a bottleneck size of 512, which selected from a pool of 32.8k/25.1k tokens. For the Triangle dataset, we used a bottleneck size of 64 from a pool of 2.0k tokens. For the relational reasoning tasks, we used a bottleneck size of 256, which selected from a pool of 14.4k tokens. Based on the bottleneck size and the patch pool size, we used 128 memory slots for the Pet and ImageNet100 datasets and 32 memory slots for the other datasets. Moreover, we trained the models on the Pet dataset for 300 epochs and on the other datasets for 100 epochs. In relational reasoning tasks, we trained all models for 100 epochs with a batch size of 64.
@@
Optimizer AdamW
Weight decay 0.01
Learning rate $1\times 10^{-4}$
Number of self-attention heads 12
Number of attention layers 2 (Small)/ 12 (Base)
Size of hidden layer 768
Size of MLP 3072
Size of memory slot 32
Number of bottleneck attention heads 8
Beta 1.0
Epochs 100 (300 for Oxford Pet)
@@
Patch size 4
Batch size 512
Number of memory slots 32
Bottleneck size 512
@@
Patch size 32
Batch size 512
Number of memory slots 32
Bottleneck size 64
Oxford Pet and ImageNet100
Patch size 16
Batch size 128
Number of memory slots 128
Bottleneck size 512
@@
Patch size 5
Batch size 64
Number of memory slots 32
Bottleneck size 256
Table 8: Hyperparameters.
Appendix D Algorithm
Algorithm 1 Global Workspace Layer
1:  Main program
2:  Input: tokens from the previous layers: $V\in\mathbb{R}^{B\times N\times E}$; learnable low-rank priors: $\gamma\in\mathbb{R}^{M\times D}$
3:  Squash layer concatenates tokens in the batch: $\Xi\in\mathbb{R}^{BN\times E}\leftarrow V\in\mathbb{R}^{B\times N\times E}$
4:  Project $\Xi$ to a latent space with a dimension $D\ll E$: $\Xi W^{K}_{i}\in\mathbb{R}^{BN\times D}$
5:  Obtain the attention scores over the projected tokens using priors $\gamma$: $\text{A}_{i}(\gamma,\Xi)=\text{softmax}(\frac{\gamma(\Xi W^{K}_{i})^{T}}{\sqrt{D}})$
6:  Tokens compete to write in memory via a bottleneck with capacity $k$: $h_{i}=\text{top-}k(\text{A}_{i})\Xi W^{V}$ (*Bottleneck Attention Balance Loss is employed for a more diverse token selection)
7:  Update priors $\gamma^{t}$ with Exponentially Weighted Moving Average: $\hat{\gamma}^{t+1}=(1-\alpha)\cdot\gamma^{t}+\alpha\cdot\text{LN}(\text{Concat}(h_{1},\dots,h_{S})W^{O})$
8:  Layer normalization: $\gamma^{t+1}=\frac{{\hat{\gamma}^{t+1}}}{{\sqrt{\sum_{{j=1}}^{M}(\hat{\gamma}^{t+1}_{j})^{2}}}}$
9:  Project $\gamma^{t+1}\in\mathbb{R}^{M\times D}$ into a dimension of $E$ as attractors within associative memory: $f_{\text{LT}}(\gamma^{t+1})\in\mathbb{R}^{M\times E}$
10:  Reconstruct $\Xi$ using attractors $f_{\text{LT}}(\gamma^{t+1})$ with a continuous Hopfield network in one inner step of energy reduction: $\hat{\Xi}^{t}=\arg\min_{\Xi^{t}}(-\text{lse}(\beta,f_{\text{LT}}(\gamma^{t+1})\Xi^{t})+\frac{1}{2}\Xi^{t}{\Xi^{t}}^{T}+\beta^{-1}\text{log}M+\frac{1}{2}(\max_{i}|f_{\text{LT}}(\gamma^{t+1}_{i})|)^{2})$
11:  Reshape the tokens into batches as the output of the Global Workspace layer: $\hat{V}^{t}\in\mathbb{R}^{B\times N\times E}\leftarrow\hat{\Xi}^{t}\in\mathbb{R}^{(B\times N)\times E}$
12:  Bottleneck Attention Balance Loss
13:  Cumulative attention loss: $\ell_{\text{importance}_{i,o}}=\sum_{j=1}^{M}\text{A}_{i,j,o}$
14:  Selected instance loss: $\ell_{\text{loads}_{i,o}}=\sum_{j=1}^{M}(\text{A}_{i,j,o}>0)$
15:  For each attention head $i$: $\ell_{\text{bottleneck}_{i}}=\frac{{\text{Var}(\{\ell_{\text{importance}_{i,o}}}\}_{o=1}^{B\times N})}{{(\frac{1}{B\times N}\sum_{o=1}^{B\times N}\ell_{\text{importance}_{i,o}}})^{2}+\epsilon}+\frac{{\text{Var}(\{\ell_{\text{loads}_{i,o}}}\}_{o=1}^{B\times N})}{{(\frac{1}{B\times N}\sum_{o=1}^{B\times N}\ell_{\text{loads}_{i,o}}})^{2}+\epsilon}$
16:  Sum the losses over all heads: $\sum_{i=1}^{S}\ell_{\text{bottleneck}_{i}}$
Appendix E Analysis of attention head operating modes
@@
We assume that the competition within the pair-wise attention is important for the model to learn meaningful representations. If such competition exists, a trained model will naturally result in sparser interactions in attention heads. Therefore, we first performed an analysis of the operating modes of different attention heads in a pretrained ViT model by measuring the number of tokens each head is attending to. We trained a ViT-Base variant model for 100 epochs from scratch for the CIFAR-10 task. Then, for each attention head, we obtained a violin plot to represent the distribution of attention sparsity for different tokens (Figure 5). The attention sparsity for a specific patch’s interactions with other tokens is computed as follows $\arg\min_{s}\sum_{j=1}^{s}A^{i,j}\geq 0.9$, where $A^{i,j}$ is the attention score allocated to the $j$th patch by the $i$th patch. The attention sparsity score is measured by the minimal number of required tokens whose attention scores add up to 0.90. For instance, there are 65 tokens for the CIFAR-10 task with patch size 4, thus there are 65 interactions for each patch to all the tokens including itself. An attention head has a higher sparsity if the median $\bar{s}$ of the required tokens is smaller, which is the number in the center of each panel. The heads in each layer are sorted according to $\bar{s}$. Note that training the model for a longer duration can result in even better convergence and higher attention sparsity. We also refer to a concurrent investigation on the Bidirectional Encoder Representations from Transformers (BERT) for results on the natural language processing (NLP) tasks [22].
@@
Figure 5: Analysis of operating modes of attention heads in the ViT-Base model. We recognize three different groups of attention heads based on their sparsity scores. Group (I) in light blue: High sparsity heads abundant in the middle layers 3-6. The vast majority of these heads only used 50% or fewer interactions. Group (II) in orange: Middle sparsity heads predominant in layers 2 and 7-10. Less than 80% of the interactions were activated. Group (III) in red: Low sparsity heads observed in high layers 11-12 and the first layer, where the most tokens were attended to. The global workspace layer will provide the inductive bias to attend to the essential tokens more effectively.
@@
Compared to NLP tasks, much less sparsity was found in ViT models. This can be attributed to the long-range dependencies of the attention mechanism were more frequently employed in handling patches of images, compared to handling tokens in text sequences. In ViT models, an input image is divided into non-overlapping patches which are then treated as separate ”tokens” and fed into the transformer model. By contrast, the sequential nature of the NLP text data allows the model to capture long-range dependencies among long-range tokens more easily. Furthermore, depending on the observation of sparsity, the bottleneck size proposed in this work can vary across layers in practice, and different datasets with varying complexities require different capacities as well. When a model has fewer layers, a constant bottleneck size can perform competitively with a well-tuned one. However, for deeper models, further investigation is needed to determine the optimal capacity size. For example, middle layers might benefit more from a smaller bottleneck compared to other layers.
Appendix F Ablation study during test time
During the test time, we included a new ‘W/O Attention’  ablation, where the multi-head cross-attention is disabled, and only the Hopfield networks are leveraged for the retrieval from the explicit memory. We evaluated ViT performance in both image classification (with AiT-Medium) and relational reasoning tasks (with AiT-Small). Table 9 showed that the multi-head cross-attention with the bottleneck is important for improving AiT’s performance during the test time.
Methods CIFAR10 (%) CIFAR100 (%) Triangle (%) Oxford Pet (%) Relational tasks (%)
AiT 84.59 60.58 99.57 30.05 76.82
W/O Attention 84.50 60.56 99.56 28.68 75.17
Table 9: Performance comparison with the W/O Attention ablation.
Appendix G Efficacy of the bottleneck attention balance loss